OpenStudy (anonymous):

Reduce the fraction to lowest terms. Do not use spaces in your answer. x^3-x^2 _______ = x^4

OpenStudy (anonymous):

\[\frac{x^3-x^2}{x^4}=\frac{x^2(x-1)}{x^4}\] is a start, then you can cancel the \(x^2\) from top and bottom